OpenStudy (anonymous):

write in matrix form: x'=y, y'=x+4 my teacher has not gone over this, I need help

OpenStudy (amistre64):

what is matrix form?

OpenStudy (amistre64):

some sort of subject that this pertains to would help out as well

OpenStudy (anonymous):

I am working with differential equations

OpenStudy (amistre64):

diffy qs, thats a start :)

OpenStudy (amistre64):

and is this the start of a problem or like midways thru?

OpenStudy (anonymous):

does x'=Ax+b help?

OpenStudy (anonymous):

sorry, full question: write each system of differential equations in matrix form, i.e. x'=Ax+b

OpenStudy (amistre64):

let me look that up to see if im familiar with it by another name ...

OpenStudy (amistre64):

this looks useful http://tutorial.math.lamar.edu/Classes/DE/SystemsDE.aspx

OpenStudy (anonymous):

yeah, its actually helping me on other questions, but not this one. But thanks. :)

OpenStudy (amistre64):

is your vector [x',y'] by chance?

OpenStudy (anonymous):

idk, a vector was not mentioned

OpenStudy (amistre64):

x'=0x+y+0 y'=x+0y+4 \[\binom{x'}{y'}=\begin{pmatrix}0&1&0\\1&0&4 \end{pmatrix}\binom{x}{y}\] maybe

OpenStudy (amistre64):

matrixes are vectors ...

OpenStudy (amistre64):

\[\binom{x'}{y'}=\begin{pmatrix}0&1\\1&0 \end{pmatrix}\binom{x}{y}+\binom{0}{4}\] maybe

OpenStudy (mr.math):

We have the system \(x'=0x+y \) \(y'=x+0y+4\). We can write this as: \({x' \choose y'}=\left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right]{x\choose y}+{0\choose 4}.\)

OpenStudy (amistre64):

yay!! thats good enough for validation to me :)

OpenStudy (anonymous):

really? i thought it was more involved than that, cool. Thanks guys!

OpenStudy (mr.math):

Although I think it can be solved without finding any eigenvalues or eigenvectors.